# Globally Rigid Convex Braced Polygons

### v1

@@ 2 Convex Braced Polygons
Let $f,g:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$ be two piecewise constant, upper-semi-continuous, monotone increasing functions, with a finite number of values on $[0,2\pi)$, and $l>0$, such that for all real $x$, $f(x+l)=f(x)+2\pi$ and $g(x+l)=g(x)+2\pi$. Then there is a fixed constant $\theta$ such that $|f(x)-g(x)-\theta|<\pi$.
Proof. Notice that by replacing, say $f(x)$ by $f(x)+c_{1}$, and $g(x)$ with $g(x)+c_{2}$ where $c_{1},c_{2}$ are any real constants, it does not change the conditions on $f$ and $g$. From the period conditions, $f(x)-g(x)$ is periodic with period $l$, as well as being upper semi-continuous. From the piecewise constant conditions, there is a minimum point $m$ and a maximum point $M$ such that for all $x$ in $[m,m+l)$ and $M$ in $[m,m+l)$,
$f(m)-g(m)\leq f(x)-g(x)\leq f(M)-g(M),$ (1)
with at least one of the inequalities being strict. This is since if both inequalities are equal for all $x$, then
$f(m)-g(m)=f(x)-g(x)=f(M)-g(M)$
for all $x$, and $f(x)=g(x+Nl)$ for all $x$ for some integer $N$, and the conclusion holds easily. So, henceforward, we will assume that one of the inequalities of (1) is strict.
Then, by the inequality above, the periodicity, and the assumptions on $f$ and $g$, we have
$0<f(M)-g(M)-(f(m)-g(m))=f(M)-f(m)-(g(M)-g(m))\leq 2\pi-0=2\pi.$
Replace $f(x)$ with $F(x)=f(x)-f(m)$, and $g(x)$ with $G(x)=g(x)-g(m)$. So inequality (1) becomes
$0\leq F(x)-G(x)\leq f(M)-f(m)-(g(M)-g(m))=F(M)-G(M)<2\pi.$
In other words, the difference between the maximum value and the minimum value of $F(x)-G(x)$ is strictly less than $2\pi$. Then let
$\theta=(F(M)-G(M))/2$
the average value of the extreme values of $F(x)-G(x)$. This is what was desired, since the mid-point in an interval is no further to every point in the interval than half the length of the interval, and $\theta$ varies continuously with $f$ and $g$. ∎
Figure 2: These are two oriented pairs of line segments joined head to tail with the difference between the directions shown.
@@
Figure 3: This shows two convex angles, i.e. oriented clockwise from $1$ to $2$, such that their average is not convex. But the $b_{2}$ bar direction has been rotated more than $\pi$, and this causes the average angle to be not convex.
Up to now, effectively we have been considering the curvature at the vertices of a polygon, as in Figure 2. Now consider the two strictly convex polygons themselves, with $n$ vertices, $(G,{\bf p})$ and $(G,{\bf q})$ with corresponding sides equal. For a point $x$ on the segment $[p_{i},p_{i+1})$, let $\theta_{i}=f(x)$, and $\phi_{i}=g(x)$ be the angle that an edge vector makes with a fixed direction in the plane so that as one proceeds around the polygons $(G,{\bf p})$ and $(G,{\bf q})$, the $i$-th angle is the direction $\theta_{i}$ of $p_{i+1}-p_{i}$ and the direction $\phi_{i}$ of $q_{i+1}-q_{i}$, respectively, both modulo $2\pi$. We make it so that both $\theta_{i}$ and $\phi_{i}$ are monotone increasing, when the orientation of the polygon is in the clockwise sense, so that $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$, $\phi_{i}<\phi_{i+1}<\phi_{i}+\pi$, and defined for all integers $i=\dots,-1,0,1,\dots$. The restriction to $\pi$ is because of the local convexity assumption on $(G,{\bf p})$ and $(G,{\bf q})$. Also for all $i$, $\theta_{i+n}=\theta_{i}+2\pi$ and $\phi_{i+n}=\phi_{i}+2\pi$, where $G$ has $n$ vertices. The maps $f$ and $g$ “unwrap” the the map from the boundary of the polygon to the circle to a similar map from the line to the line, as a function of the curve length.
The condition for strict local convexity for $(G,{\bf p})$ is that $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$, and, similarly, for $(G,{\bf q})$, it is that $\phi_{i}<\phi_{i+1}<\phi_{i}+\pi$. Even when both polygons are convex at the given corresponding vertices, Figure 3 shows that the angle for the average may not be convex for the given orientation. This is because the $b_{2}$ bar has been rotated clockwise more than $\pi$ to $c_{2}$. Local convexity can only be destroyed when one of the edges in ${\bf p}+{\bf q}$ has the opposite orientation, and this only happens when the angles between corresponding edges is $\pi$ or greater. (Note that, for computing the average configuration, it is a translate of the average of any translates of $p$ and $q$. This allows more freedom in drawing averages.) We apply Lemma 2.1 to this situation, and when $l$ is the perimeter of the polygon, and we define a piecewise constant function for both $(G,{\bf p})$ and $(G,{\bf q})$ to obtain monotone increasing piecewise constant functions.
Proof of Theorem 2.1: We have shown that when there is a non-trivial infinitesimal flex, there are non-congruent strictly convex bar frameworks $(G,{\bf p})$ and $(G,{\bf q})$. Here we assume that there are two non-congruent strictly convex braced bar polygons $(G,{\bf p})$ and $(G,{\bf q})$ with corresponding bars of the same length, both oriented clockwise, or both oriented counter-clockwise. Thus, there is piecewise-linear map from the boundary polygon of $(G,{\bf p})$ to the boundary polygon of $(G,{\bf q})$, where each boundary is of length $l$. For a point in the relative interior of each edge $p_{i},p_{i+1}$ and at $p_{i}$, define $f(x)$ to be the oriented angle in the direction $(p_{i+1}-p_{i})/||p_{i+1}-p_{i}||$ as an angle on the unit circle. Notice that $f$ is upper semi-continuous. Similarly, define $g$ for $(G,{\bf q})$. (Notice that $f$ and $g$ are a sort of piecewise Gaus map for the convex polygons.)
By Lemma 2.1 one can add an angle $\theta$ to the function $f$, which is equivalent to rotating $(G,{\bf p})$ by $\theta$ so that the average $(G,({\bf p}+{\bf q})/2)$ is strictly convex. Then ${\bf p}-{\bf q}$ is a non-trivial infinitesimal flex of $(G,({\bf p}+{\bf q})/2)$. ∎
@@
Even when one strictly convex braced polygon $(G,{\bf p})$ has one of its edges parallel to the corresponding edge of $(G,{\bf q})$, then the average $(G,({\bf p}+{\bf q})/2)$ may not be convex as seen in Figure 4.
Figure 4: On the top figures, this shows two strictly convex braced polygons with corresponding sides the same length, and one pair $\{1,2\}$ of corresponding sides that are parallel in the same direction, but their average is not a convex polygon. The bottom figures show the same two strictly convex braced polygons but with one rotated so that their average is convex. The dotted line segment between corresponding vertices have the average vertices at their midpoint.
@@
Figure 5 shows two corners where one pair $b_{1},c_{1}$ have equal length with one orientation, but the other pair $b_{2},c_{2}$ do not have equal length, but have the same orientation. However, the average pair $(b_{1}+c_{1})/2,(b_{2}+c_{2})/2$ have the opposite orientation. So the sign of the angle between the successive bars can change when corresponding bars do not have the same length.
Figure 5: These are two corners, where not all corresponding sides are equal, and the average configuration has the opposite orientation
@@
Figure 6 shows two bar frameworks that have corresponding bars of the same length. The one on the right is a strictly convex braced polygons that is infinitesimally rigid in all strictly convex realizations and so it is convexly rigid by Theorem 2.1. On the other hand, even though the graph of these polygons are globally rigid in the plane for all generic configurations, due to the configuration on the left, these configurations are necessarily not globally rigid in the plane and the one on the left is necessarily not strictly convex.
@@ 3 Braced Polygons with Proper Stresses
In this section, we shall strengthen Theorem 2.1 by restricting our attention to special families of braced polygons. Our discussion is motivated from the result by Geleji and Jordan
[Geleji and Jordán(2013)], which characterizes the global rigidity of all strictly convex realizations when the underlying graph is a circuit in the generic rigidity matroid. We shall develop a novel analysis of the rigidity of braced polygons by focusing on minimal 3-connectivity.
@@ 3.1 Proper Stresses and Super Stability
Our tool to prove global rigidity (or even universal rigidity) is super stability based on self-stresses.
@@
For a configuration $(G,{\bf p})$, a stress is a function that maps the set of edges to real numbers, $\omega:E\to\mathbb{R}$. The stress is in equilibrium if $\sum_{j}\omega_{ij}(\mathbf{p}_{i}-\mathbf{p}_{j})=\mathbf{0}$ for every vertex $i$. A braced polygon is properly stressed if it has an equilibrium stress that is positive on the boundary polygon and negative on the interior edges. The stress matrix $\Omega$ of a stress $\omega$ is an $n\times n$ symmetric matrix defined as $\Omega_{ij}=-\omega_{ji}$ for each edge $e_{ij}$, $\Omega_{ij}=0$ if $i\not=j$ and there is no edge between $v_{i}$ and $v_{j}$, and that each row sums to 00.
@@
A basic theorem in [Connelly and Guest(2022), Theorem 5.14.1] states that super stable frameworks are universally rigid. The reverse is not true. See Figure 21E in the plane, and see [Connelly and Gortler(2015)] for the general situation.
@@
Suppose a strictly convex braced polygon $(G,{\mathbf{p}})$ is properly stressed. Then it is super stable for the corresponding stress matrix.
@@
Proof. At each vertex consider the three vectors, centered at the vertex; the sum of the interior edges weighted by their non-positive stresses, and the two boundary edges weighted by their stresses. The equilibrium condition implies that the sum of these three vectors is 00. If just one of these vectors is 00, it implies that the other two are negatives of each other, and this contradicts the strict convexity at that vertex. Similarly, if two of the vectors are 00, so is the third and they are all 00. Then the local strict convexity at the vertex implies that when all three are non-zero, the boundary vertex stresses are both positive by the independence of the three vectors.
@@ 3.2 Minimally 3-connected Braced Polygons with Proper Stresses
In view of Theorem 3.1, the next natural question would be to identify strictly convex braced polygons having proper stresses. We have a concrete answer for the family of minimally 3-connected braced polygons, whose formal definition is given as follows.
@@
A braced polygon is minimally $k$-connected if it is $k$-connected, and removing any brace (interior edge) would cause it to be no longer $k$ vertex connected.
@@
Every 3-connected braced polygon $P$ has a strictly convex realization that is properly stressed. In addition, if $B$ is a set of braces such that $P-B$ is 3-connected, then there exists a strictly convex realization such that $P-B$ is properly stressed and all braces in $B$ have positive stresses.
Proving Theorem 3.3 requires several additional ingredients that will be proved individually. First, we show there is an inductive procedure to produce any 3-connected braced polygon. Next, we show that if a convex but not strictly convex braced polygon is properly stressed, then there exists a nearby strictly convex configuration with desired stress by minimizing a certain energy function.
@@
Every 3-connected braced polygon $P$ with at least 4 vertices can be constructed by adding one brace at a time by the following operations starting from $Q=K_{4}$:
@@
The graph $Q$ is 3-connected in every step. Furthermore, if $P$ is minimally 3 vertex connected, then it is constructed by only using the first 2 operations.
@@
There exists a differentiable non-negative strictly monotonically increasing real function $f$ defined on $[0,\infty)$ such that given $a,\omega,m,\delta>0$, $f(a)=\frac{1}{m}$, $f^{\prime}(a)=\omega$, and $f(a+\delta)>1$. Similarly, there exists a differentiable positive strictly monotonically decreasing function $g$ defined on $(0,\infty)$ such that given $a,m,\delta>0$ and $\omega<0$, $g(a)=\frac{1}{m}$, $g^{\prime}(a)=\omega$, and $g(a-\delta)>1$.
@@
To see that $\mathbf{q}$ is strictly convex, all vertices are in a small neighborhood of $\mathbf{p}$, therefore, every vertex that was strictly convex stays so with a sufficiently small $\delta$. For the newly added vertices, they have degree 3 with positive stress on the boundary and negative stress on a brace, so they must be strictly convex. This concludes the last part of Theorem 3.3.
By the same energy minimization argument, if an edge is not necessary for the 3 vertex connectivity, then it can have either positive or negative stress. We can pick $\epsilon$ to be either positive or negative in this case and a small $\delta$. A point in the small neighborhood of a strictly convex polygon must also be strictly convex. ∎
@@ 3.3 Minimal 3-connectivity and circuits
Minimally 3-connected braced polygons also enjoy the following special property.
@@
Let $(G,{\mathbf{p}})$ be a minimally 3-connected strictly convex polygon and $\omega$ be a non-zero equilibrium stress of $(G,{\mathbf{p}})$ that is non-positive on the interior braces. Then $\omega$ is a proper stress and the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension one.
Proof. By Proposition 3.2, $\omega$ is positive on the boundary edges. Suppose $\omega$ is zero on some interior brace. Let $Z$ be the set of interior braces which have zero stress in $\omega$. Then, $\omega$ is a proper stress of $(G-Z,{\mathbf{p}})$. This implies from Theorem 3.1 that $G-Z$ is 3-connected. Since $Z$ contains at least one brace, this contradicts the minimal 3-connectivity of $G$. Thus $\omega$ is a proper stress.
Suppose that the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension at least two. Let $\omega_{1}$ be a proper stress and $\omega_{2}$ be an arbitrary equilibrium stress that is not a scalar multiple of $\omega_{1}$. Consider $\omega_{t}:=\omega_{1}+t\omega_{2}$ for a parameter $t$ starting at $t=0$, and increase $t$ continuously. Let $\bar{t}$ be the smallest $t$ such that $\omega_{t}(e)=0$ holds for some edge $e$. Then the existence of $\omega_{\bar{t}}$ contradicts the former part of the statement. ∎
Generic rigidity circuits are one of the central classes of graphs in the context of global rigidity.
@@
A graph $G$, with with $n$ vertices and $m$ edges, is called a (planar) circuit if $m=2n-2$ and for every subgraph on $k$ vertices such that $2\leq k<n$, the number of edges is smaller than or equal to $32k-$ 3.
@@
Let $G$ be a minimally 3-connected polygon graph with $n$ vertices and $m$ edges. If $m\geq 2n-2$, then $G$ is a circuit. Moreover, every strictly convex framework $(G,{\mathbf{p}})$ is properly stressed, and $m=n-2$.
@@
Suppose $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid. Since $(G,{\mathbf{p}}_{0})$ is infinitesimally rigid, we have $t>0$. We may assume that $(G,{\mathbf{p}}_{t^{\prime}})$ is infinitesimally rigid and for any $t^{\prime}\in[0,t)$ and is properly stressed (by the same argument as above). We consider a sequence $\{{\mathbf{p}}_{t_{i}}\}_{i=1,2.\dots}$ of point configurations that converges to ${\mathbf{p}}_{t}$. Let $\omega_{t_{i}}$ be a proper stress of $(G,{\mathbf{p}}_{t_{i}})$ having unit norm. By the compactness of the unit sphere, a subsequence of $\{\omega_{t_{i}}\}_{i=1,2,\dots}$ converges to a nonzero $\omega$, which is an equilibrium stress of $(G,{\mathbf{p}}_{t})$. Since $\omega$ is non-positive on the interior braces, Lemma 3.1 implies that $\omega$ is a proper stress. However, since $G$ is a circuit and $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid, the space of equilibrium stressed of $(G,{\mathbf{p}}_{t})$ is at least two, contradicting Lemma 3.1. Thus also $m=n-2$ because $G$ is a circuit. This completes the proof. ∎

### v2

@@ 2 Convex Braced Polygons
Let $f,g:\mathbb{R}\rightarrow\mathbb{R}$ be monotone step functions. Suppose that $(f,g)$ is periodic in the sense that, for some positive numbers $l$ and $C$, $f(x+l)=f(x)+C$ and $g(x+l)=g(x)+C$ for all $x\in\mathbb{R}$. Suppose further that $f(x)\in[0,C)$ and $g(x)\in[0,C)$ for all $x\in[0,l)$. Then there is a constant $\alpha$ such that $|f(x)-g(x)-\alpha|<\frac{C}{2}$ for all $x\in\mathbb{R}$.
Proof. By the assumption on $f$ and $g$, $f-g$ is a periodic piecewise constant function with period $l$. Then $f-g$ takes only finitely many values so has a minimizer $m$ and a maximizer $M$. By the periodicity of $f-g$, we may assume $M\in[m,m+l)$. Then,
$0\leq f(M)-f(m)\leq C\mbox{ and }0\leq g(M)-g(m)\leq C.$ (2)
Since $g(m)<C$ and $g(l)\geq C$, if $g(m)=g(M)$, then $M<l$ holds and hence $f(M)<C$ holds. In particular, by $f(m)\geq 0$, $f(M)-f(m)=C$ and $g(M)=g(m)$ cannot occur simultaneously. Therefore, we have
00 0 $\leq f(x)-g(x)-(f(m)-g(m))$
$\leq f(M)-g(M)-(f(m)-g(m))$
$=f(M)-f(m)-(g(M)-g(m))$
$<C-0=C,$
where the first inequality follows from the definition of $m$, the second inequality follows from the definition of $M$, and the fourth strict inequality follows from (2) and the fact that $f(M)-f(m)=C$ and $g(M)=g(m)$ cannot occur simultaneously. Let $F(x)=f(x)-f(m)$ and $G(x)=g(x)-g(m)$. The above inequality gives
$0\leq F(x)-G(x)\leq F(M)-G(M)<C$ (3)
for any $x\in\mathbb{R}$. We set
$\alpha=f(m)-g(m)+\frac{F(M)-G(M)}{2}.$
Then, for any $x$, $f(x)-g(x)-\alpha=F(x)-G(x)-\frac{F(M)-G(M)}{2}$. Hence, by (3), $|f(x)-g(x)-\alpha|\leq|\frac{F(M)-G(M)}{2}|<\frac{C}{2}$. ∎
@@
Figure 3: This shows two convex angles such that their average is not convex.
We are now ready to prove Theorem 2.1.
Proof of Theorem 2.1: We have already seen that, if $G$ has a strictly convex realisation with a non-trivial infinitesimal flex, then $G$ has two non-congruent strictly convex realisations in which corresponding bars have the same edge lengths (by Proposition 2.2).
To see the converse direction, we assume that there are two non-congruent strictly convex braced polygons $(G,{\bf p})$ and $(G,{\bf q})$ with corresponding bars of the same length, both oriented counter-clockwise. For $x\in\mathbb{R}$, let $p_{x}$ be the point on the boundary polygon of $(G,{\mathbf{p}})$ whose distance from $p_{1}$ along the polygon in the counter-clockwise direction is equal to $x$. Consider $\theta_{i}$ as defined before Lemma 2.1 based on ${\mathbf{p}}$, and define $f:\mathbb{R}\rightarrow\mathbb{R}$ by $f(x)=\theta_{i}$ for $x\in\mathbb{R}$ with $p_{x}\in[p_{i},p_{i+1})$. Similarly, $g:\mathbb{R}\rightarrow\mathbb{R}$ is defined based on ${\mathbf{q}}$. Since $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}})$ have corresponding bar lengths the same, $f$ and $g$ are monotone step functions and they are periodic with $f(x+l)=f(x)+2\pi$ and $g(x+l)=g(x)+2\pi$ for any $x\in\mathbb{R}$, where $l$ is the circumference of the boundary polygons of $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}})$. By Lemma 2.1, there is an $\alpha$ such that $|f(x)-g(x)-\alpha|<\pi$. This in turn implies that, if $(G,{\mathbf{q}}^{\prime})$ is obtained by rotating $(G,{\mathbf{q}})$ by $\alpha$, (1) holds between $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}}^{\prime})$, and $(G,\frac{{\mathbf{p}}+{\mathbf{q}}^{\prime}}{2})$ forms a braced strictly convex polygon. By Proposition 2.3, $(G,\frac{{\mathbf{p}}+{\mathbf{q}}^{\prime}}{2})$ is not infinitesimally rigid. This completes the proof. ∎
@@
Even when one strictly convex braced polygon $(G,{\bf p})$ has one of its edges parallel to the corresponding edge of $(G,{\bf q})$, the average $(G,({\bf p}+{\bf q})/2)$ may not be convex as seen in Figure 4.
Figure 4: The top figure shows two strictly convex braced polygons with corresponding sides the same length, and with one pair $\{1,2\}$ of corresponding sides that are parallel and have the same direction, but their average is not a convex polygon. The bottom figure shows the same two strictly convex braced polygons but with one rotated so that their average is convex. The dotted line segment between corresponding vertices have the average vertices at their midpoint.
@@
Figure 5 shows two frameworks consisting of two bars $b_{1},b_{2}$ and $c_{1},c_{2}$. Bars $b_{1},c_{1}$ have equal length but the other pair $b_{2},c_{2}$ do not have equal length. Both $b_{1},b_{2}$ and $c_{1},c_{2}$ have the same orientation. However, the average pair $(b_{1}+c_{1})/2,(b_{2}+c_{2})/2$ has the opposite orientation. So the sign of the angle between successive bars can change when corresponding bars do not have the same length.
Figure 5: Two frameworks, where not all corresponding pairs of bars have equal length, and the average configuration has the opposite orientation
@@
Figure 6 shows two non-congruent frameworks that have corresponding bars of the same length. Hence neither framework is globally rigid. The one on the right is a strictly convex braced polygon that is infinitesimally rigid in all strictly convex realizations and so is convexly rigid by Theorem 2.1. Thus the left framework is necessarily not strictly convex. The underlying graph of these frameworks is globally rigid in the plane at all generic configurations.
@@ 3 Braced Polygons with Proper Stresses
In this section, we shall strengthen Theorem 2.1 by restricting our attention to special families of braced polygons. Our discussion is motivated by a result by Geleji and Jordan
[Geleji and Jordán(2013)], which characterizes the global rigidity of all strictly convex plane realizations of a braced polygon graph which is a circuit in the generic 2-dimensional rigidity matroid. We shall develop a novel analysis of the rigidity of braced polygon graphs by focusing on minimal 3-connectivity.
@@ 3.1 Proper Stresses and Super Stability
Our tool to prove global rigidity (or even universal rigidity) is super stability, which is based on the concept of an equilibrium stress.
@@
For a configuration $(G,{\bf p})$, a stress is a function $\omega:E\to\mathbb{R}$ that maps the set of edges of $G$ to the real numbers. The stress is in equilibrium if $\sum_{j}\omega_{ij}(\mathbf{p}_{i}-\mathbf{p}_{j})=\mathbf{0}$ for every vertex $i$, taking $\omega_{ij}=0$ when $ij\not\in E$. The stress matrix $\Omega$ of a stress $\omega$ is an $n\times n$ symmetric matrix defined as $\Omega_{ij}=-\omega_{ji}$ for each edge $e_{ij}$, $\Omega_{ij}=0$ if $i\not=j$ and there is no edge between $v_{i}$ and $v_{j}$, and that each row sums to 00.
We say that an equilibrium stress for a convex braced polygon is proper if it is positive on the boundary polygon and negative on the interior edges, and that the convex braced polygon is properly stressed if it has a proper stress.
@@
A basic theorem in [Connelly and Guest(2022), Theorem 5.14.1] states that super stable frameworks are universally rigid. The reverse is not true. See Figure 20 in the plane, and see [Connelly and Gortler(2015)] for the general situation.
@@
Suppose a strictly convex braced polygon $(G,{\mathbf{p}})$ has a proper stress. Then it is super stable for the corresponding stress matrix.
@@
Proof. At each vertex, consider the three vectors, centered at the vertex; the sum of the interior edges weighted by their non-positive stresses, and the two boundary edges weighted by their stresses. The equilibrium condition implies that the sum of these three vectors is 00. If just one of these vectors is 00, it implies that the other two are negatives of each other, and this contradicts the strict convexity at that vertex. Similarly, if two of the vectors are 00, so is the third and they are all 00. Then the local strict convexity at the vertex implies that when all three are non-zero, the boundary vertex stresses are both positive by the independence of the three vectors.
@@ 3.2 Minimally 3-connected Braced Polygons with Proper Stresses
In view of Theorem 3.1, the next natural question is to identify the strictly convex braced polygons which have a proper stress. We give a positive answer for the family of strictly convex realisations of minimally 3-connected braced polygon graphs, whose formal definition is given as follows.
@@
A braced polygon graph is minimally $k$-connected if it is $k$-connected, and removing any brace (interior edge) would cause it to be no longer $k$-connected.
@@
Every 3-connected braced polygon graph $P$ has a strictly convex realization that is properly stressed. In addition, if $B$ is a set of braces such that $P-B$ is 3-connected, then there exists a strictly convex realization of $P$ such that $P-B$ is properly stressed and all braces in $B$ have positive stresses.
Proving Theorem 3.3 requires several additional ingredients that will be proved individually. First, we show there is an inductive procedure to produce any 3-connected braced polygon graph. Next, we show that if a convex but not strictly convex braced polygon is properly stressed, then there exists a nearby strictly convex configuration with a proper stress by minimizing a certain energy function.
@@
Every 3-connected braced polygon graph $P$ with at least 4 vertices can be constructed from $K_{4}$ by adding one brace at a time using the following operations:
@@
In addition, 3-connectivity is preserved at every step of the above construction and, if $P$ is minimally 3-connected, then it can be constructed by only using the first 2 operations.
@@
There exists a differentiable non-negative strictly monotonically increasing real function $f$ defined on $[0,\infty)$ such that given $a,m,\delta,\omega>0$, we have $f(a)=\frac{1}{m}$, $f^{\prime}(a)=\omega$, and $f(a+\delta)>1$. Similarly, there exists a differentiable positive strictly monotonically decreasing function $g$ defined on $(0,\infty)$ such that given $a,m,\delta>0$ and $\omega<0$, we have $g(a)=\frac{1}{m}$, $g^{\prime}(a)=\omega$, and $g(a-\delta)>1$.
@@
To see that $\mathbf{q}$ is strictly convex, all vertices are in a small neighborhood of $\mathbf{p}$, therefore, every vertex that was strictly convex stays so with a sufficiently small $\delta$. For the newly added vertices, they have degree 3 with positive stress on the boundary and negative stress on a brace, so they must be strictly convex.
By the same energy minimization argument, if an edge is not necessary for the 3 vertex connectivity, then it can have either positive or negative stress. We can pick $\epsilon$ to be either positive or negative in this case and a small $\delta$. A point in the small neighborhood of a strictly convex polygon must also be strictly convex. This concludes the last part of Theorem 3.3. ∎
@@ 3.3 Minimal 3-connectivity and circuits
Minimally 3-connected braced polygon graphs enjoy the following special property.
@@
Let $G$ be a minimally 3-connected braced polygon graph, $(G,{\bf p})$ be a strictly convex realisation of $G$ and $\omega$ be a non-zero equilibrium stress of $(G,{\mathbf{p}})$ that is non-positive on the interior braces. Then $\omega$ is a proper stress and the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension one.
Proof. By Proposition 3.2, $\omega$ is positive on the boundary edges. Suppose $\omega$ is zero on some interior brace. Let $Z$ be the set of interior braces which have zero stress in $\omega$. Then, $\omega$ is a proper stress of $(G-Z,{\mathbf{p}})$. This implies from Theorem 3.1 that $G-Z$ is 3-connected. Since $Z$ is non-empty, this contradicts the minimal 3-connectivity of $G$. Thus $\omega$ is a proper stress.
Suppose that the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension at least two. Let $\omega_{1}$ be a proper stress and $\omega_{2}$ be an arbitrary equilibrium stress that is not a scalar multiple of $\omega_{1}$. We may assume that $\omega_{1}(e)$ and $\omega_{2}(e)$ have different signs on some edge $e$ (by taking the negative of $\omega_{2}$ if necessary).
Consider $\omega_{t}:=\omega_{1}+t\omega_{2}$ for a parameter $t$ starting at $t=0$, and increase $t$ continuously. Let $\bar{t}$ be the smallest $t$ such that $\omega_{t}(e)=0$ holds for some edge $e$. Since $\omega_{\bar{t}}$ is non-positive on the interior braces, the existence of $\omega_{\bar{t}}$ contradicts the former part of the statement. ∎
Graphs whose edge sets are circuits in the generic $2$-dimensional rigidity matroid are important in the study of global rigidity in $\mathbb{R}^{2}$. We next consider braced polygon graphs which are generic rigidity circuits.
@@
A graph $G$, with with $n$ vertices and $m$ edges, is called a generic rigidity circuit if $m=2n-2$ and for every subgraph on $k$ vertices such that $2\leq k<n$, the number of edges is smaller than or equal to $32k-$ 3. A braced polygon graph which is a generic rigidity circuit is simply called a braced polygonal circuit.
@@
Let $G$ be a minimally 3-connected braced polygon graph with $n$ vertices and $m$ edges. If $m\geq 2n-2$, then $G$ is a generic rigidity circuit and every strictly convex braced polygon $(G,{\mathbf{p}})$ is properly stressed.
@@
Suppose $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid for some $t$. Since $(G,{\mathbf{p}}_{0})$ is infinitesimally rigid, we have $t>0$. We may assume that $(G,{\mathbf{p}}_{t^{\prime}})$ is infinitesimally rigid for any $t^{\prime}\in[0,t)$ and is properly stressed (by the same argument as above). We consider a sequence $\{{\mathbf{p}}_{t_{i}}\}_{i=1,2.\dots}$ of point configurations that converges to ${\mathbf{p}}_{t}$. Let $\omega_{t_{i}}$ be a proper stress of $(G,{\mathbf{p}}_{t_{i}})$ having unit norm. By the compactness of the unit sphere, a subsequence of $\{\omega_{t_{i}}\}_{i=1,2,\dots}$ converges to a nonzero $\omega$, which is an equilibrium stress of $(G,{\mathbf{p}}_{t})$. Since $\omega$ is non-positive on the interior braces, Lemma 3.1 implies that $\omega$ is a proper stress. However, since $G$ is a circuit and $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid, the space of equilibrium stressed of $(G,{\mathbf{p}}_{t})$ is at least two, contradicting Lemma 3.1. This completes the proof. ∎